If C is a subspace of F, prove that dim(C) dim(C)= n .

(C is a binary linear code with length n and dimension k, C^\perp is the dual code of C)
Answer #1

le cis a subspace of the show linke) + dim (et) = n. Since oua subspace of 5n, then S i ... z o be the basis for a Then din cSo nullsface of (AT) =n-ki Ince rank (A) = rank (AT), from eg we get sant (A) ant rank(A) + nullity CAT) = dim (Col(A)) + dim